# CTPU-PTC-24-02

### v1

@@
${}^{a}$ Department of Physics, Keio University, Yokohama, 223-8522, Japan
${}^{b}$ Center for Theoretical Physics of the Universe, Institute for Basic Science (IBS), Daejeon 34051, Korea
${}^{c}$ Department of Physics, Hokkaido University, Sapporo 060-0810, Japan
@@ 1 Introduction
The finite modular symmetry has been intensively studied to explain the flavor structure of the Standard Model (SM) quarks and leptons [1, 2, 3, 4, 5, 6, 7, 8, 9]. (See for reviews [10, 11].) Under the symmetry, a Yukawa coupling constant behaves as the so-called modular form which is a holomorphic function of the modulus $\tau$. Interestingly, the finite modular symmetries $\Gamma_{N}$, $N\in\mathbb{N}$, are isomorphic to the non-Abelian discrete symmetries [12], which can explain the flavor structure of the SM fermions [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such modular dependence of the Yukawa coupling to a modulus was found in heterotic orbifold models [23, 24, 25, 26, 27, 28] and magnetized D-brane models [29, 30, 31, 32, 33, 34, 35] in superstring theory. Moreover, the finite modular symmetry can explain the hierarchies of the quark and lepton masses and mixing if the modulus is stabilized near one of the fixed points where the residual symmetry remains unbroken [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
In the models with finite modular flavor symmetries, the value of modulus $\tau$ plays an essential role to fit to the data and is often treated as a free parameter. Whereas in a UV theory, its value should be fixed dynamically through moduli stabilization. There are studies on the moduli stabilization with the three-form fluxes [48, 49], the negative power of modular form [50, 51] 111 The positive power of the trivial singlet is also studied in Ref. [51]. See also Ref. [52]. , and the general superpotential invariant under the modular symmetry $SL(2,\mathbb{Z})$ [53, 54, 55, 56, 57], utilizing the Klein $j$ function [58]. The latter two mechanisms will need non-perturbative dynamics to generate those potentials. Recently, effect of the radiative correction to the tree-level stabilization is discussed in Ref. [59].
In this work, we point out that the modulus can be stabilized only by the Coleman-Weinberg (CW) potential generated by couplings with matter fields. This mechanism provides a perturbative way to stabilize a modulus through the supersymmetry breaking. For illustration, we shall consider a simple model with vector-like quarks which transforms as non-trivial singlets under the finite modular flavor symmetry. The resultant potential has a global minimum at $\mathrm{Im}\,\tau\gg 1$, where the residual $\mathbb{Z}^{T}_{N}$ symmetry associated with the $T$ transformation, $\tau\to\tau+1$, is unbroken. That can lead to a large hierarchy suppressing instanton-like correction, i.e., $e^{-2\pi{\mathrm{Im}\,\tau}}\ll 1$. An interesting feature of this mechanism is that the $\mathrm{Re}\,\tau$ direction can be identified as the QCD axion due to the accidental Peccei-Quinn (PQ) symmetry originated from the $\mathbb{Z}^{T}_{N}$ residual symmetry. We shall discuss the condition that the axion solution to the strong CP problem is not spoiled due to the CW potential, i.e. the axion quality is good enough.
@@ 3 Radiative moduli stabilization
We consider a simple supersymmetric model with the following Kähler potneital and superpotnetial:
@@
$V_{\mathrm{CW}}=\frac{1}{32\pi^{2}}\sum_{i}$ $\left[\left(m_{i}^{2}+m_{Q_{i}}^{2}(\tau)\right)^{2}\left(\log\left(\frac{m_{i}^{2}+m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right.$ (3.2)
$\hskip 85.35826pt\left.-(m_{Q_{i}}^{2}(\tau))^{2}\left(\log\left(\frac{m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right],$
@@ 3.1 Simplified analysis
where the second parenthesis of Eq. (3.6) is vanishing. Here, $\mathcal{W}(z)$ is the Lambert function satisfying $\mathcal{W}(z)e^{\mathcal{W}(z)}=z$ which has two real values for $-e^{-1}<z<0$, and the larger one can be a solution where our approximation $\left|{q}\right|\ll 1$ is valid. Taking $\mu=\tilde{M}_{Q}$ and $p=2\pi/3~{}$, the location of the minimum is $x_{0}=3.9,~{}7.9,~{}12$ and $16$ for $k=6,8,10$ and $12$, respectively. $x_{0}$ increases slightly by decreasing $\mu$. Thus we find that there is the minimum at $\mathrm{Im}\,\tau\gg 1$ in the CW potential. Note that this minimum exists only for $t>0$ where the modular form $Y^{({k})}_{1_{t}}$ is a non-trivial singlet. For the CW potential from the trivial-singlet modular form, the potential increases as $\mathrm{Im}\,\tau$ increases in contrast to that from the non-trivial one, and thus there is no minimum at $\mathrm{Im}\,\tau\gg 1$ 333 The potentials discussed in Refs. [53, 54, 55, 56, 57, 58] have a similar behavior, that is, $V\to\infty$ for $\mathrm{Im}\,\tau\to\infty$. See also the potential in Ref. [52]. Therefore the dominant CW potential should be generated from the non-trivial singlet one for the existence of the minimum at $\mathrm{Im}\,\tau\gg 1$, so that the potential tends to approach to zero for a sufficiently large ${\rm Im}\tau$.
@@ 3.2 Numerical analysis
We shall study the potential Eq. (3.2) numerically without relying on the $q$-expansion in Eq. (3.4). For illustration, we consider $\Gamma_{3}\simeq A_{4}$ as the finite modular group, and two pairs of vector-like quarks whose superpotential is given by
@@
The modular weights are chosen to $k=12$ so that the strong CP problem is solved with keeping the modulus mass heavy, as discussed in the next section 555 Since $1_{1}$ and $1_{2}$ can be in the $2^{\prime}$ representation under $A_{4}^{\prime}$, the two fields could be embedded into a doublet. . Here, we introduce the two vector-like pairs, so that the mixed anomaly of $A_{4}$ with the QCD is canceled [61, 62]. The singlets are trivial under the $S$-transformation, and transform as $Y^{({k})}_{1_{t}}\to w^{t}Y^{({k})}_{1_{t}}$ under the $T$-transformation. Hence, in our setup, $\det(\rho(T))=w^{1+2}=1$ and the anomaly is canceled. The anomaly associated with the modular weight can be canceled by the Green-Schwarz mechanism with an another gauge coupling modulus or by adding $\log Y^{({k})}_{1_{0}}(\tau)$ to the gauge kinetic function with a certain coefficient, where $Y^{({k})}_{1_{0}}(\tau)$ is a trivial singlet modular form and could be explained by threshold corrections from heavy modes [63].
@@ 4 Axion solution and its quality
We have seen that the CW potential Eq. (3.2) induced by the superpotential Eq. (3.8) is very flat in the $\mathrm{Re}\,\tau$ direction. The stabilization mechanism can realize the axion solution to the strong CP problem by assuming that the matter fields are vector-like quarks as in the KSVZ axion model [64, 65]. The effective $\theta$-angle $\overline{\theta}$ in the QCD is given by
@@
where $\alpha$ is a real transformation parameter. Hence the PQ charge of $\overline{Q}_{t}Q_{t}$ is $-t$, and the $U(1)_{\rm PQ}$ is anomalous. This symmetry is ensured from the residual $\mathbb{Z}^{T}_{3}$ symmetry corresponding to $\alpha=2\pi$, and the explicit breaking effects of the $U(1)_{\mathrm{PQ}}$ are from $q=e^{2\pi i\tau}$ which is invariant under $\mathbb{Z}^{T}_{3}$ but violates $U(1)_{\mathrm{PQ}}$ 666 The corrections to the axion from $\log Y^{({k})}_{1_{0}}$ in the gauge kinetic function which may exist to cancel the weight anomaly will also appear at $\mathcal{O}\left({q}\right)$. . Therefore, $\mathrm{Re}\,\tau$ can be identified as the QCD axion. Since the kinetic term of the modulus is given by 777 We write explicitly the reduced Planck mass $M_{p}$ in this section.
@@
$\left|{\Delta\theta}\right|\simeq 2\times 10^{-10}\times\sin\theta_{0}\left(\frac{\left|{\Omega}\right|}{10^{4}}\right)\left(\frac{m_{0}}{10^{7}~{}\mathrm{GeV}}\right)^{2}\left(\frac{M_{Q}}{M_{p}}\right)^{2}\left(\frac{x_{0}}{28}\right)^{8}\left(\frac{\epsilon}{10^{-12}}\right)^{7}.$ (4.8)
@@
in the $A_{4}$ model. Thus the modulus mass decreases as the axion quality is higher by $\epsilon\to 0$. For $\Delta\theta/\sin\theta_{0}\sim 10^{-10}$, $\epsilon\lesssim 10^{-12}$ is required to avoid the moduli problem 999 The lighter modulus would not be a problem if there is the thermal inflation [74, 75, 76], large Hubble induced mass during the inflation exists [77, 78], and/or an inflation scale is lower than the modulus mass. .
The mass of the lighter vector-like quark $Q_{2}$, whose $T$ charge is $2$, is given by
@@
$s_{Q_{i}}\sim$ $\ \frac{{\lambda}_{i}x_{0}^{7}Y^{({14})}_{1_{1}}{\langle{H_{d}}\rangle}}{M_{Q}x_{0}^{6}Y^{({12})}_{1_{i}}}\sim\frac{{\lambda}_{i}x_{0}{\langle{H_{d}}\rangle}}{M_{Q}\epsilon^{i-1}}$ (4.14)
$\sim$ $\ 10^{-3}\times{\lambda}_{2}\left(\frac{x}{28}\right)\left(\frac{{\langle{H_{d}}\rangle}}{100~{}\mathrm{GeV}}\right)\left(\frac{M_{p}}{M_{Q}}\right)\left(\frac{10^{-12}}{\epsilon}\right),$
@@ 5 Summary and discussions
In this work, we point out that the modulus of the finite modular symmetry can be stabilized by the Coleman-Weinberg (CW) potential. For illustration, we studied the model with $\Gamma_{3}\simeq A_{4}$ symmetry and two vector-like quark pairs which transform as non-trivial singlets, namely $1_{1}$ and $1_{2}$. This is the minimal possibility to cancel the mixed QCD anomaly of the finite modular symmetry, but to induce that of the $U(1)_{\mathrm{PQ}}$ symmetry. The CW potential has the global minimum at $\mathrm{Im}\,\tau\gg 1$ if the corresponding modular form is non-trivial singlet under $A_{4}$, where the residual symmetry $\mathbb{Z}^{T}_{3}$ remains unbroken. Since the potential along the $\mathrm{Re}\,\tau$ direction has extremely flat due to this residual symmetry, we can regard $\mathrm{Re}\,\tau$ as the QCD axion to solve the strong CP problem. The accidental $U(1)_{\mathrm{PQ}}$ arises due to the $\mathbb{Z}^{T}_{3}$ symmetry and hence the PQ breaking effects are controlled by $\Gamma_{3}$ which is a different situation argued in Ref. [84]. We examined the condition to ensure the quality of this finite modular axion, and correlate with the masses of the modulus $X\sim\mathrm{Im}\,\tau$ and vector-like quarks. Interestingly, the modulus $X$ and the lighter quark $Q_{2}$ are expected to be $\mathcal{O}\left({\mathrm{TeV}}\right)$ scale for $\Delta\theta<10^{-10}$ and thus could be probed by cosmology and collider experiments. This mechanism can be generalized to other modular forms as long as the CW potential is dominated by that with non-zero $T$-charge and the potential converges to zero for $\mathrm{Im}\,\tau\gg 1$.
@@
The modular $A_{4}$ symmetry in our scenario can not be used to explain the hierarchical structure of the SM fermions, as studied in Refs. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], since the hierarchy parameter $\epsilon$ is $\mathcal{O}\left({10^{-12}}\right)$ and is too small. Such a tiny $\epsilon$ is required due to the correlation of the axion quality $\Delta\theta$ with the modulus mass in Eq. (4.10). This relation can be relaxed if the $\tau$-independent angle $\theta_{0}$ is small or the level $N$ is large. In the latter case, there exist a small hierarchy $|e^{2\pi i\tau/N}|$ in the SM sector and a large one $|e^{2\pi i\tau}|$ in the vector-like quark sector.101010 In general, for example, if one could break the periodicity $\tau\sim\tau+1$ to $\tau\sim\tau+1/n$ ($n>1$) in the vector-like $(Q,\bar{Q})$ sector with keeping the periodicity $\tau\sim\tau+1$ in the SM fermion sector, then modular forms may appear $Y^{k}_{r}(\tau)$ in the SM fermion sector and $Y^{k^{\prime}}_{r^{\prime}}(n\tau)$ in the $(Q,\bar{Q})$ sector, and one could realize both a small hierarchy $|e^{2\pi i\tau}|$ in the SM sector and a large one $|e^{2\pi in\tau}|$ in the vector-like quark pairs. . Thus, we could construct a model which explains the SM fermion hierarchies by the modulus stabilized by the mechanism proposed in this paper. An explicit model building is subject of our future work.

### v2

@@
a Department of Physics, Keio University, Yokohama, 223-8522, Japan
b Center for Theoretical Physics of the Universe, Institute for Basic Science (IBS), Daejeon 34051, Korea
c Department of Physics, Hokkaido University, Sapporo 060-0810, Japan
@@ 1 Introduction
The finite modular symmetry has been intensively studied to explain the flavor structure of the Standard Model (SM) quarks and leptons [1, 2, 3, 4, 5, 6, 7, 8, 9]. (See for reviews [10, 11].) Under the symmetry, a Yukawa coupling constant behaves as a so-called modular form which is a holomorphic function of the modulus $\tau$. Interestingly, the finite modular symmetries $\Gamma_{N}$, $N\in\mathbb{N}$, are isomorphic to the non-Abelian discrete symmetries [12], which can explain the flavor structure of the SM fermions [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such modular dependence of the Yukawa coupling to a modulus was found in heterotic orbifold models [23, 24, 25, 26, 27, 28] and magnetized D-brane models [29, 30, 31, 32, 33, 34, 35] in superstring theory. Moreover, the finite modular symmetry can explain the hierarchies of the quark and lepton masses and mixing if the modulus is stabilized near one of the fixed points where the residual symmetry remains unbroken [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47].
In the models with finite modular flavor symmetries, the value of modulus $\tau$ plays an essential role to fit to the data and is often treated as a free parameter. Whereas in a UV theory, its value should be fixed dynamically through modulus stabilization. There are studies on the moduli stabilization with the three-form fluxes [48, 49], the negative power of modular form [50, 51] 111 The positive power of the trivial singlet is also studied in Ref. [51]. See also Ref. [52]. , and the general superpotential invariant under the modular symmetry $SL(2,\mathbb{Z})$ [53, 54, 55, 56, 57], utilizing the Klein $j$ function [58]. The latter two mechanisms will need non-perturbative dynamics to generate those potentials. Recently, effect of the radiative correction to the tree-level stabilization is discussed in Ref. [59].
In this work, we point out that the modulus can be stabilized only by the Coleman-Weinberg (CW) potential generated by couplings with matter fields. This mechanism provides a perturbative way to stabilize a modulus through the supersymmetry breaking. For illustration, we shall consider a simple model with vector-like quarks which transform as non-trivial singlets under the finite modular flavor symmetry. The resultant potential has a global minimum at $\mathrm{Im}\,\tau\gg 1$, where the residual $\mathbb{Z}^{T}_{N}$ symmetry associated with the $T$ transformation, $\tau\to\tau+1$, is unbroken. That can lead to a large hierarchy suppressing instanton-like correction, i.e., $e^{-2\pi{\mathrm{Im}\,\tau}}\ll 1$. An interesting feature of this mechanism is that the $\mathrm{Re}\,\tau$ direction can be identified as the QCD axion due to the accidental Peccei-Quinn (PQ) symmetry originated from the $\mathbb{Z}^{T}_{N}$ residual symmetry. We shall discuss the condition that the axion solution to the strong CP problem is not spoiled due to the CW potential, i.e. the axion quality is good enough.
@@ 3 Radiative moduli stabilization
We consider a simple supersymmetric model with the following Kähler potential and superpotential:
@@
$V_{\mathrm{CW}}=\frac{1}{32\pi^{2}}\sum_{i}$ $\left[\left(m_{i}^{2}+m_{Q_{i}}^{2}(\tau)\right)^{2}\left(\log\left(\frac{m_{i}^{2}+m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right.$
$\hskip 85.35826pt\left.-(m_{Q_{i}}^{2}(\tau))^{2}\left(\log\left(\frac{m_{Q_{i}}^{2}(\tau)}{\mu^{2}}\right)-\frac{3}{2}\right)\right],$ (3.2)
@@ 3.1 Simplified analysis
where the second parenthesis of Eq. (3.6) is vanishing. Here, $\mathcal{W}(z)$ is the Lambert function satisfying $\mathcal{W}(z)e^{\mathcal{W}(z)}=z$ which has two real values for $-e^{-1}<z<0$, and the larger one can be a solution where our approximation $\left|{q}\right|\ll 1$ is valid. Taking $\mu=\tilde{M}_{Q}$ and $p=2\pi/3~{}$, the location of the minimum is $x_{0}=3.9,~{}7.9,~{}12$ and $16$ for $k=6,8,10$ and $12$, respectively. $x_{0}$ increases slightly by decreasing $\mu$. Thus we find that there is the minimum at $\mathrm{Im}\,\tau\gg 1$ in the CW potential. Note that this minimum exists only for $t\neq 0$ where the modular form $Y^{({k})}_{1_{t}}$ is a non-trivial singlet. For the CW potential from the trivial-singlet modular form, the potential increases as $\mathrm{Im}\,\tau$ increases in contrast to that from the non-trivial one, and thus there is no minimum at $\mathrm{Im}\,\tau\gg 1$ 333 The potentials discussed in Refs. [53, 54, 55, 56, 57, 58] have a similar behavior, that is, $V\to\infty$ for $\mathrm{Im}\,\tau\to\infty$.. See also the potential in Ref. [52]. Therefore the dominant CW potential should be generated from the non-trivial singlet one for the existence of the minimum at $\mathrm{Im}\,\tau\gg 1$, so that the potential tends to approach to zero for a sufficiently large ${\rm Im}\tau$.
@@ 3.2 Numerical analysis
We shall study the potential Eq. (3) numerically without relying on the $q$-expansion in Eq. (3.4). For illustration, we consider $\Gamma_{3}\simeq A_{4}$ as the finite modular group, and two pairs of vector-like quarks whose superpotential is given by
@@
The modular weights are chosen to $k=12$ so that the strong CP problem is solved with keeping the modulus mass heavy, as discussed in the next section 555 Since $1_{1}$ and $1_{2}$ can be in the $2^{\prime}$ representation under $A_{4}^{\prime}$, the two modular forms could be embedded into a doublet. . Here, we introduce the two vector-like pairs, so that the mixed anomaly of $A_{4}$ with the QCD is canceled [61, 62]. The singlets are trivial under the $S$-transformation, and transform as $Y^{({k})}_{1_{t}}\to w^{t}Y^{({k})}_{1_{t}}$ under the $T$-transformation. Hence, in our setup, $\det(\rho(T))=w^{1+2}=1$ and the anomaly is canceled. The anomaly associated with the modular weight can be canceled by the Green-Schwarz mechanism with an another gauge coupling modulus or by adding $\log Y^{({k})}_{1_{0}}(\tau)$ to the gauge kinetic function with a certain coefficient, where $Y^{({k})}_{1_{0}}(\tau)$ is a trivial singlet modular form and could be explained by threshold corrections from heavy modes [63].
@@ 4 Axion solution and its quality
We have seen that the CW potential Eq. (3) induced by the superpotential Eq. (3.8) is very flat in the $\mathrm{Re}\,\tau$ direction. The stabilization mechanism can realize the axion solution to the strong CP problem by assuming that the matter fields are vector-like quarks as in the KSVZ axion model [64, 65]. The effective $\theta$-angle $\overline{\theta}$ in the QCD is given by
@@
where $\alpha$ is a real transformation parameter. Hence, the PQ charge of $\overline{Q}_{t}Q_{t}$ is $-t/3$, and the $U(1)_{\rm PQ}$ is anomalous. This symmetry is ensured from the residual $\mathbb{Z}^{T}_{3}$ symmetry corresponding to $\alpha=2\pi$, and the explicit breaking effects of the $U(1)_{\mathrm{PQ}}$ are from $q=e^{2\pi i\tau}$ which is invariant under $\mathbb{Z}^{T}_{3}$ but violates $U(1)_{\mathrm{PQ}}$ 666 The corrections to the axion from $\log Y^{({k})}_{1_{0}}$ in the gauge kinetic function which may exist to cancel the weight anomaly will also appear at $\mathcal{O}\left({q}\right)$. . Therefore, $\mathrm{Re}\,\tau$ can be identified as the QCD axion. Since the kinetic term of the modulus is given by 777 We write explicitly the reduced Planck mass $M_{p}$ in this section.
@@
$\left|{\Delta\theta}\right|\simeq 2\times 10^{-10}\times\sin\theta_{0}\left(\frac{\left|{\Omega}\right|}{10^{4}}\right)\left(\frac{m_{0}}{10^{7}~{}\mathrm{GeV}}\right)^{2}\left(\frac{M_{Q}}{M_{p}}\right)^{2}\left(\frac{x_{0}}{28}\right)^{12}\left(\frac{\epsilon}{10^{-12}}\right)^{7}.$ (4.8)
@@
in the $A_{4}$ model. Thus the modulus mass decreases as the axion quality is higher by $\epsilon\to 0$. For $\Delta\theta/\sin\theta_{0}\sim 10^{-10}$, $\epsilon\lesssim 10^{-12}$ is required to avoid the moduli problem 999 The lighter modulus would not be a problem if there is the thermal inflation [74, 75, 76], large Hubble induced mass during the inflation exists [77, 78], and/or an inflation scale is lower than the modulus mass. The amount of relativistic axions produced from modulus decays, which may affect to $\Delta N_{\mathrm{eff}}$ is discussed in Refs.  [79, 80, 81]. .
The mass of the lighter vector-like quark $Q_{2}$, whose charge under $\mathbb{Z}^{T}_{3}$ is $2$, is given by
@@
$s_{Q_{i}}\sim$ $\ \frac{{\lambda}_{i}x_{0}^{7}Y^{({14})}_{1_{1}}{\langle{H_{d}}\rangle}}{M_{Q}x_{0}^{6}Y^{({12})}_{1_{i}}}\sim\frac{{\lambda}_{i}x_{0}{\langle{H_{d}}\rangle}}{M_{Q}\epsilon^{i-1}}$
$\sim$ $\ 10^{-3}\times{\lambda}_{2}\left(\frac{x}{28}\right)\left(\frac{{\langle{H_{d}}\rangle}}{100~{}\mathrm{GeV}}\right)\left(\frac{M_{p}}{M_{Q}}\right)\left(\frac{10^{-12}}{\epsilon}\right),$ (4.14)
@@ 5 Summary and discussions
In this work, we point out that the modulus of the finite modular symmetry can be stabilized by the Coleman-Weinberg (CW) potential. For illustration, we studied the model with $\Gamma_{3}\simeq A_{4}$ symmetry and two vector-like quark pairs which transform as non-trivial singlets, namely $1_{1}$ and $1_{2}$. This is the minimal possibility to cancel the mixed QCD anomaly of the finite modular symmetry, but to induce that of the $U(1)_{\mathrm{PQ}}$ symmetry. The CW potential has the global minimum at $\mathrm{Im}\,\tau\gg 1$ if the corresponding modular form is non-trivial singlet under $A_{4}$, where the residual symmetry $\mathbb{Z}^{T}_{3}$ remains unbroken. Since the potential along the $\mathrm{Re}\,\tau$ direction is extremely flat due to this residual symmetry, we can regard $\mathrm{Re}\,\tau$ as the QCD axion to solve the strong CP problem. The accidental $U(1)_{\mathrm{PQ}}$ arises due to the $\mathbb{Z}^{T}_{3}$ symmetry and hence the PQ breaking effects are controlled by $\Gamma_{3}$ which is a different situation argued in Ref. [87]. We examined the condition to ensure the quality of this finite modular axion, and correlate with the masses of the modulus $X\sim\mathrm{Im}\,\tau$ and vector-like quarks. Interestingly, the modulus $X$ and the lighter quark $Q_{2}$ are expected to be $\mathcal{O}\left({\mathrm{TeV}}\right)$ scale for $\Delta\theta<10^{-10}$ and thus could be probed by cosmology and collider experiments. This mechanism can be generalized to other modular forms as long as the CW potential is dominated by that with non-zero $\mathbb{Z}^{T}_{3}$-charge and the potential converges to zero for $\mathrm{Im}\,\tau\gg 1$.
@@
The modular $A_{4}$ symmetry in our scenario can not be used to explain the hierarchical structure of the SM fermions, as studied in Refs. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], since the hierarchy parameter $\epsilon$ is $\mathcal{O}\left({10^{-12}}\right)$ and is too small. Such a tiny $\epsilon$ is required due to the correlation of the axion quality $\Delta\theta$ with the modulus mass in Eq. (4.10). This relation can be relaxed if the $\tau$-independent angle $\theta_{0}$ is small or the level $N$ is large. In the latter case, there exists a small hierarchy $|e^{2\pi i\tau/N}|$ in the SM sector and a large one $|e^{2\pi i\tau}|$ in the vector-like quark sector.101010 In general, for example, if one could break the periodicity $\tau\sim\tau+1$ to $\tau\sim\tau+1/n$ ($n>1$) in the vector-like $(Q,\bar{Q})$ sector with keeping the periodicity $\tau\sim\tau+1$ in the SM fermion sector, then modular forms may appear $Y^{k}_{r}(\tau)$ in the SM fermion sector and $Y^{k^{\prime}}_{r^{\prime}}(n\tau)$ in the $(Q,\bar{Q})$ sector, and one could realize both a small hierarchy $|e^{2\pi i\tau}|$ in the SM sector and a large one $|e^{2\pi in\tau}|$ in the vector-like quark pairs. . Thus, we could construct a model which explains the SM fermion hierarchies by the modulus stabilized by the mechanism proposed in this paper. An explicit model building is subject of our future work.